# MaLA-500: Massive Language Adaptation of Large Language Models

### v1

@@ 3.1 Comparison across LLMs
In our comprehensive analysis of contributing factors, a significant Pearson correlation of 0.53 is observed between the effect of vocabulary extension and the performance gain. This observation holds true for languages with both non-Latin scripts, such as Kannada (kan_Knda), Malayalam (mal_Mlym), and Tigrinya (tir_Ethi), as well as Latin scripts, such as Igbo (ibo_Latn) and Yoruba (yor_Latn). Furthermore, we note that the corpus size of a language exhibits a weak correlation of 0.13 with its performance gain. In contrast, the corpus size of the language family to which a language belongs demonstrates a moderate correlation of 0.40.
@@ 3.2 Effect of Number of Shots
Fig. 2 illustrates the relationship between accuracy and the number of in-context examples (i.e., shots). As the number of in-context shots increases, there is a corresponding rise in accuracy. Notably, with just 1-shot, accuracy exhibits randomness at 26.08%, indicating 1-shot provides limited information for task learning. The transition from 1 shot to 2 shots/3 shots results in a significant improvement, with performances boosted by 20.97% and 28.02%, respectively. This highlights the effectiveness of increasing the number of shots. MaLA-500 achieves its peak performance at approximately 60% accuracy with 6-10 in-context shots. This may be attributed to the multi-class nature of the SIB-200 dataset, necessitating more shots for learning intricate input-label mappings.
In Fig. 3, a more nuanced portrayal of results aligns with the observations made in Fig. 2. In the realm of 1-shot in-context learning, approximately 40 languages exhibit erratic results. As the number of shots increases, there is a reduction in the number of languages achieving low accuracy (20-40%), coupled with a growing cohort achieving high accuracy (60-80%). Notably, in the case of 10-shot in-context learning, 101 (57%) languages surpass the threshold of 60% accuracy.
@@
Figure 2: In-context learning macro-average accuracy (%) on SIB-200 with different number of shots using MaLA-500.
Figure 3: Detailed results of in-context learning on SIB-200 using MaLA-500. X-axis: the number of languages in different accuracy ranges (%). Y-axis: number of shots.
@@ 4.1 Multilingual Language Models
Language model development has endeavored to broaden the scope of pertaining languages to address multilingual scenarios. Pretrained multilingual models have been able to accommodate up to a hundred or more languages. Noteworthy examples include mBERT Devlin et al. (2019), which supports 104 languages, XLM-R (Conneau et al., 2020) covering 100 languages, mBART (Liu et al., 2020) designed for 25 languages, mT5 (Xue et al., 2021) spanning 101 languages, XGLM (Lin et al., 2021) across 30 languages, GPT-3 covering 118 languages (93% English), mGPT (Shliazhko et al., 2022) accommodating 60 languages, and BLOOM (Scao et al., 2022) supporting 46 languages and 13 programming languages.
Surprisingly, two recent multilingual language models have surpassed the conventional limit by supporting more than 400 languages. Glot500-m (ImaniGooghari et al., 2023) spans 511 languages through vocabulary extension and continued training based on XLM-R. SERENGETI (Adebara et al., 2022) goes even further by supporting 517 African languages and language varieties, written in five different scripts, employing models inspired by both ELECTRA (Clark et al., 2020) and XLM-R. MADLAD (Kudugunta et al., 2023) covers 419 languages and trains an 8B language model from scratch with an adapted UL2 objective (Tay et al., 2022). Our work is concurrent with the MADLAD400 language model. We distinguish it from: 1) language coverage. Our work covered more than 500 languages, a number comparable to that of encoder-only models and surpassing MADLAD-400 by an additional 100 languages. 2) training methods. We consider continual training to benefit from the learned knowledge of the original models. 3) model architecture. We adopt an open model architecture, i.e., LLaMA, while MADLAD uses UL2.
@@ 4.2 Language Adaptation
Before the advent of LLMs, diverse approaches are employed to adapt small-scale multilingual language models to new languages. These methods include using adapters (Pfeiffer et al., 2020; Üstün et al., 2020; Pfeiffer et al., 2020; Nguyen et al., 2021; Faisal and Anastasopoulos, 2022; Yong et al., 2022), vocabulary extension and substitution (Chau et al., 2020; Wang et al., 2020; Müller et al., 2020, 2021; Pfeiffer et al., 2021; Chen et al., 2023b; Downey et al., 2023), leveraging monolingual corpora (Ebrahimi and Kann, 2021), and utilizing bilingual lexicons (Wang et al., 2022).
While language models have been scaled up significantly, their coverage is limited to a specific set of languages. To address this constraint, various methods have been proposed to expand the applicability of these Large Language Models (LLMs) across a broader range of languages, catering to both general-purpose tasks and specific applications like machine translation. These methods also involve vocabulary extension (Cui et al., 2023), continued pretraining and instruction-tuning (Yong et al., 2022; Cui et al., 2023; Chen et al., 2023a), and parallel corpora exploitation (Cahyawijaya et al., 2023; Yang et al., 2023; Zhu et al., 2023; Xu et al., 2023). Despite these efforts, massive language adaptation of LLMs for general-purpose tasks across diverse languages remains an area yet to be thoroughly explored.
5 Conclusion
We present a pioneering effort in massive language adaptation on LLMs, focusing on extending LLaMA 7B to our model, MaLA-500. This adaptation involves vocabulary extension and continued pretraining with LoRA. Our approach leads to MaLA-500 achieving state-of-the-art in-context learning capabilities, as demonstrated on the task of SIB-200. We release the model weights publicly to facilitate future research. This work marks a substantial advancement in applying LLMs to a diverse range of languages.
Limitations
Data
We do not collect more high-resource data and instead use only Glot500-c, which does not put full effort on collecting data for high-resource languages. In addition, the flores200-based SIB-200 evaluation set is included in the training data since Glot500-c includes flores200, but the classification labels are not provided.
Model Size
Our study limits the maximum model size to 10B parameters. In our future work, we plan to expand to a larger model.
Ethical Implications
@@ Acknowledgements
This work was funded by the European Research Council (grants #740516 and #758969), EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), and the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070350 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number 10052546]. The authors wish to acknowledge CSC – IT Center for Science, Finland, for generous computational resources. Shaoxiong Ji and Peiqin Lin acknowledge travel support from ELISE (GA no 951847).
@@ Appendix A Languages
The list of languages of Glot500-c used to train MaLA-500 with the number of available sentences and language family information for each language is available in Tables 2,  3 and 4.
Lang |Sent| Family Lang |Sent| Family Lang |Sent| Family
@@
Lang |Sent| Family Lang |Sent| Family Lang |Sent| Family
@@
Lang |Sent| Family Lang |Sent| Family Lang |Sent| Family

### v2

@@ 3.3 Comparison across Languages
In our comprehensive analysis of contributing factors on SIB200, we note that the corpus size of a language exhibits a weak correlation of 0.13 with its performance gain. In contrast, the corpus size of the language family to which a language belongs demonstrates a moderate correlation of 0.40. A moderately high Pearson correlation of 0.53 is observed between the effect of vocabulary extension, i.e., the reduction in segmentation length, and the performance gain. This observation holds true for languages with both non-Latin scripts, such as Kannada (kan_Knda), Malayalam (mal_Mlym), and Tigrinya (tir_Ethi), as well as Latin scripts, such as Igbo (ibo_Latn) and Yoruba (yor_Latn). It demonstrates the effectiveness of vocabulary extension.
@@ 3.4 Effect of Number of Shots
Figure 4 illustrates the relationship between accuracy and the number of in-context examples (i.e., shots) on SIB200. As the number of in-context shots increases, there is a corresponding rise in accuracy. Notably, with just 1-shot, accuracy exhibits randomness at 30.88%, indicating 1-shot provides limited information for task learning. The transition from 1 shot to 2 shots/3 shots results in a notable improvement, with performances boosted by 19.83% and 26.14%, respectively. This highlights the effectiveness of increasing the number of shots. MaLA-500 achieves its peak performance at approximately 65% accuracy with 6-10 in-context shots. This may be attributed to the multi-class nature of the SIB200 dataset, necessitating more shots for learning intricate input-label mappings.
Figure 4: In-context learning macro-average accuracy (%) on SIB200 with different number of shots using MaLA-500.
In Figure 5, a more nuanced portrayal of results aligns with the observations made in Figure 4. In the realm of 1-shot in-context learning, approximately 50 languages exhibit erratic results. As the number of shots increases, there is a reduction in the number of languages achieving low accuracy (25-50%), coupled with a growing cohort achieving high accuracy (75-100%).
Figure 5: Detailed results of in-context learning on SIB200 using MaLA-500. X-axis: the number of languages in different accuracy ranges (%). Y-axis: number of shots.
@@ 4.1 Multilingual Language Models
Language model development has endeavored to broaden the scope of pretraining languages to address multilingual scenarios. Pretrained multilingual models have been able to accommodate up to a hundred or more languages. Noteworthy examples include mBERT Devlin et al. (2019), which supports 104 languages, XLM-R (Conneau et al., 2020) covering 100 languages, mBART (Liu et al., 2020) designed for 25 languages, mT5 (Xue et al., 2021) spanning 101 languages, XGLM (Lin et al., 2021) across 30 languages, GPT-3 covering 118 languages (93% English), mGPT (Shliazhko et al., 2022) accommodating 60 languages, and BLOOM (Scao et al., 2022) supporting 46 languages and 13 programming languages.
Surprisingly, two recent multilingual language models have surpassed the conventional limit by supporting more than 400 languages. Glot500-m (ImaniGooghari et al., 2023) spans 511 languages through vocabulary extension and continued training based on XLM-R. SERENGETI (Adebara et al., 2022) goes even further by supporting 517 African languages and language varieties, written in five different scripts, employing models inspired by both ELECTRA (Clark et al., 2020) and XLM-R. MADLAD (Kudugunta et al., 2023) covers 419 languages and trains an 8B language model from scratch with an adapted UL2 objective (Tay et al., 2022). Our work is concurrent with the MADLAD-400 language model. We distinguish it by: 1) language coverage. Our work covered more than 500 languages, a number comparable to that of encoder-only models and surpassing MADLAD-400 by an additional 100 languages. 2) training methods. We consider continual training to benefit from the learned knowledge of the original models. 3) model architecture. We adopt an open model architecture, i.e., LLaMA, while MADLAD uses decoder-only T5 architecture, which has not been supported by the HuggingFace ecosystem at the time of writing, thus leading to additional difficulty in usage.
@@ 4.2 Language Adaptation
Before the advent of LLMs, diverse approaches are employed to adapt small-scale multilingual language models to new languages. These methods include using adapters (Pfeiffer et al., 2020; Üstün et al., 2020; Pfeiffer et al., 2020; Nguyen et al., 2021; Faisal & Anastasopoulos, 2022; Yong et al., 2022), vocabulary extension and substitution (Chau et al., 2020; Wang et al., 2020; Müller et al., 2020; 2021; Pfeiffer et al., 2021; Chen et al., 2023; Downey et al., 2023), leveraging monolingual corpora (Ebrahimi & Kann, 2021; Alabi et al., 2022), and utilizing bilingual lexicons (Wang et al., 2022).
While language models have been scaled up notably, their coverage is limited to a specific set of languages. To address this constraint, various methods have been proposed to expand the applicability of these large language models across a broader range of languages, catering to both general-purpose tasks and specific applications like machine translation. These methods also involve vocabulary extension (Cui et al., 2023), continued pretraining and instruction-tuning (Yong et al., 2022; Cui et al., 2023; Chen et al., 2024; Zhao et al., 2024), and parallel corpora exploitation (Cahyawijaya et al., 2023; Yang et al., 2023; Zhu et al., 2023; Xu et al., 2023). Despite these efforts, massive language adaptation of LLMs for general-purpose tasks across diverse languages, e.g., covering many languages families and more than one hundred languages, remains an area yet to be thoroughly explored.
5 Conclusion and Future Work
We present a pioneering effort in massive language adaptation on LLMs, focusing on extending LLaMA 7B to our model, MaLA-500. This adaptation involves vocabulary extension and continued pretraining with LoRA. Our approach leads to MaLA-500 achieving state-of-the-art in-context learning capabilities, as demonstrated on the benchmarks of SIB200 and Taxi1500. We release the training scripts and model weights publicly to facilitate future research. This work marks a substantial advancement in applying LLMs to a diverse range of languages.
Our future work will focus on further improving the model capacity, for example, on machine translation across many language pairs. Alves et al. (2023) showed that LLMs (LLaMA-7B and LLaMA-13B) exhibited poor performance even on English-centric high-resource language pairs in some cases. Translation with LLMs on low-resource languages is more challenging. The LLaMA-7B model performed poorly in our preliminary experiments. Besides, our pretraining corpus does not intentionally include bilingual texts, and our MaLA-500 model is not instruction-tuned with translation data. We leave the inclusion of bilingual text during continual pretraining, instruction fine-tuning with translation data, and the evaluation on machine translation as future works.
Ethical Statement
@@ 5 Conclusion and Future Work
Reproducibility Statement
We make the following efforts to ensure reproducible research. We release the model weights (https://huggingface.co/MaLA-LM) and codes for training and evaluation (https://github.com/MaLA-LM/mala-500). We use publicly available evaluation benchmarks which can be obtained freely or by request. The results are reproducible with our released model weights and evaluation scripts,
@@ Acknowledgements
We thank José Pombal for constructive suggestions on training. This work is funded by The European Research Council (grants #740516, #771113 and #758969), EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), and the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070350 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant #10052546]. The authors wish to acknowledge CSC – IT Center for Science, Finland, for generous computational resources on the Mahti supercomputer and LUMI supercomputer through the LUMI extreme scale access (MOOMIN and LumiNMT). Shaoxiong Ji and Peiqin Lin acknowledge travel support from ELISE (GA no 951847).
@@ Appendix A Languages
The list of languages of Glot500-c used to train MaLA-500 with the number of available sentences and language family information for each language is available in Tables 5,  6 and 7.
Lang $\|Sent\|$ Family Lang $\|Sent\|$ Family Lang $\|Sent\|$ Family
@@
Lang $\|Sent\|$ Family Lang $\|Sent\|$ Family Lang $\|Sent\|$ Family
@@
Lang $\|Sent\|$ Family Lang $\|Sent\|$ Family Lang $\|Sent\|$ Family